OpenStudy (anonymous):

Integrate: (-3x^2)/(x^3 +2)^2 without using log

OpenStudy (anonymous):

substitute u = x^3

OpenStudy (anonymous):

du = 3x^2 dx

OpenStudy (anonymous):

substitude u=x^3+2 that will be more helpful

OpenStudy (anonymous):

sorry substitute u = x^3+2

OpenStudy (anonymous):

yes, thats what I meant i didnt type in the +2

OpenStudy (anonymous):

\[\int\limits (-3x^2)dx/(x^3+2)^2=\int\limits -du/u^2=1/u+c=1/(x^3+2)+c\]

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

you most welcome...